# AIDE: Task-Specific Fine Tuning with Attribute Guided MultI-Hop Data Expansion

### v1

@@ Abstract
Fine-tuning large language models (LLMs) for specific tasks requires high-quality, diverse training data relevant to the task. Recent research has leveraged LLMs to synthesize training data, but existing approaches either depend on large seed datasets or struggle to ensure both task relevance and data diversity in the generated outputs. To address these challenges, we propose AIDE, a novel data synthesis framework that uses a multi-hop process to expand $10$ seed data points while ensuring diversity and task relevance. AIDE extracts the main topic and key knowledge attributes from the seed data to guide the synthesis process. In each subsequent hop, it extracts the topic and attributes from the newly generated data and continues guided synthesis. This process repeats for a total of $K$ hops. To prevent irrelevant data generation as the hop depth increases, AIDE incorporates a residual connection mechanism and uses self-reflection to improve data quality. Our empirical results demonstrate that fine-tuning Mistral-7B, Llama-3.1-8B and Llama-3.2-3B with AIDE  achieves more than $10\%$ accuracy improvements over the base models across $13$ tasks from $5$ different benchmarks, while outperforming the models fine-tuned with state-of-the-art data synthesis methods like Evol-Instruct, DataTune and Prompt2Model.
@@
AIDE: Task-Specific Fine Tuning with Attribute Guided
MultI-Hop Data Expansion
Jiayu Li1,3,thanks: Work was done during an internship at Amazon Web Services., Xuan Zhu2, Fang Liu2, Yanjun Qi2,3 1Syracuse University, 2AWS Bedrock Science 3Correspondence: jli221@data.syr.edu, yanjunqi@amazon.com
@@ 1 Introduction
To overcome the limitation, an approach from Wei et al. (2022); Xu et al. (2022) samples task-specific training data from public NLP datasets, but the sampling often covers limited information. Another category of recent methods leverages the capabilities of LLMs to automatically generate large-scale synthetic data, enabling the training of advanced models in specific task domains. For example, Prompt2Model (Viswanathan et al., 2023) and DataTune (Gandhi et al., 2024) rely on several candidate datasets to synthesize task-specific data for fine-tuning LLMs. However, the methods either require a large set of seed data ready for rewriting or their synthetic data suffer from task relevance and diversity issues since the methods do not enforce sufficient control over the synthesis process.
@@
To address these challenges, we propose AIDE (Attribute guided multI-hop Data Expansion), a novel data synthesis framework that generates abundant training data from a small set of seed inputs. Our framework focuses on ensuring that the synthetic data maintains high task relevance, data diversity, and quality for the specific given tasks. AIDE uses LLMs as key players via a multi-hop synthesis process. Each hop in AIDE begins by extracting the main topic and important knowledge attributes from a seed sample using a LLM. This builds up knowledge triplets and then AIDE traverse a triplet consisting of the topic, a relationship, and an attribute to synthesize new data points. In the next hop, each newly generated data point becomes a seed, and the process repeats until reaching a depth of $K$ hops. This multi-hop mechanism allows for recursive data synthesis along all paths of a process tree, enabling the generation of large-volume data from just a few seeds. Extracted attributes act as control nodes in the multi-hop tree, ensuring the generated data points remain relevant to the target task. We also introduce personas as new key attributes, enhancing the generation of diverse data. As the depth of the recursive synthesis increases, the relevance of the synthetic data may diminish. To address this, we propose a residual connection mechanism to reduce irrelevance. In addition, we adapt self-reflection techniques to ensure the correctness of synthetic data, thus improving the quality of generated samples. Figure 1 illustrates the workflow of AIDE.
@@
To validate AIDE, we select three base models (Mistral-7B, Llama-3.1-8B and Llama-3.2-3B). We evaluate the performance of the base models fine-tuned with synthetic data from AIDE. We compare the results to the performance of the base models, as well as to the performance of the models fine-tuned with human-curated (gold) data and synthetic data from state-of-the-art (SOTA) data synthesis methods. Our evaluations cover diverse tasks from well-known benchmarks, including BIG-Bench bench authors (2023), MMLU Hendrycks et al. (2021), ARC-Challenge Clark et al. (2018), GSM8K Cobbe et al. (2021), and TruthfulQA Lin et al. (2022). For comparison, we include SOTA data synthesis methods such as Evol-Instruct Xu et al. (2024), DataTune Gandhi et al. (2024), and Prompt2Model Viswanathan et al. (2023). In summary, our main contributions are as follows:
@@
  • In zero-shot prompting, Mistral-7B fine-tuned with synthetic data from AIDE achieves average improvements of over 20%, 3% and 17% across tasks, compared to the base Mistral-7B, as well as Mistral-7B fine-tuned with gold training data and SOTA data synthesis methods. Additionally, AIDE enhances the performance of Llama-3.1-8B and Llama-3.2-3B, delivering average improvements of about 16% and 15% across tasks, respectively.
@@ 2 Related Work
Data synthesis for fine-tuning LLMs targets two primary problems. The first is open-domain generation, which synthesizes data across a wide range of topics and complexity levels. The second is task-specific generation, where synthetic data is tailored to a particular task. One can use the synthetic data in fine-tuning LLMs through techniques, such as instruction tuning, preference tuning, and their variations. This paper focuses on synthesizing training data for instruction tuning to enhance the performance of LLMs for specific tasks.
@@
3 Proposed Method: Attribute Guided Multi-Hop Data Expansion
Figure 1: Overview of the workflow of AIDE. $X^{(0)}_{i}$ denotes the $i$-th task-related seed data point. AIDE includes four steps. (1) a LLM extractor extracts a topic $t$, knowledge attributes $a_{1}$ and $a_{2}$ with relationships $r_{1}$ and $r_{2}$ of a data point. (2) During the multi-hop synthesis at the depth of hop $j$, a LLM acts as a synthesizer with task demonstrations ${\mathcal{D}}_{T}$ to generate data $X^{(j)}_{1}$, $X^{(j)}_{2}$ and $X^{(j)}_{3}$ along paths of synthesis with a predefined operation $Op$ (i.e., adding constraints). (3) To enhance the diversity of synthesis, we expand attributes by retrieving a persona $p_{1}$ from a persona hub with $t$. (4) A LLM grader with self-reflection ensures the quality of synthetic data. Finally, a LLM as an annotator generates the label of synthetic data. We describe the technical details of AIDE in Section 3.
3.1 Multi-Hop Synthesis
To synthesize abundant data, we propose a multi-hop synthesis approach, with an example illustrated in Figure 2. The Figure demonstrate the multi-hop synthesis process, which the seed data $X^{(0)}_{i}$ is used to synthesize its $1$-hop neighbors $X^{(1)}_{1}$ and $X^{(1)}_{2}$ during the $1$-hop synthesis. Similarly, each $1$-hop neighbor can be applied to generate $2$-hop neighbors of $X^{(0)}_{i}$. For each input data $X^{(0)}_{i}$ where $1\leq i\leq n$, we recursively synthesis data using the same pattern until reaching the depth of $K$.
Definition 3.1 (Multi-hop synthesis).
Figure 3: An example of unfolded multi-hop synthesis when $K=2$.
@@ 3.1 Multi-Hop Synthesis
Table 1: The $1$-hop synthesis in Figure 3 uses an input data point $X^{(0)}_{i}$ to generate a representation of the data point ${\mathcal{A}}^{(0)}_{i}$ with triplets $<t_{1},r_{1},a_{1}>$ and $<t_{1},r_{2},a_{2}>$. We retrieve the persona $P_{1}$ according to $t_{1}$. Through the triplets, task demonstrations ${\mathcal{D}}_{T}$, the persona $p_{1}$ and the predefined operation $Op$, we synthesize $X^{(1)}_{1}$, $X^{(1)}_{2}$ and $X^{(1)}_{3}$ by combining $X^{(0)}_{i}$ with its corresponding task category and related examples.
Figure 3 illustrates an example of unfolded multi-hop synthesis. In this example, we set $K=2$. $X^{(0)}_{i}$ is one of the seed data point and $X^{(1)}=\{X^{(1)}_{1},X^{(1)}_{2},...,X^{(1)}_{m_{1}}\}$ represents synthetic data from $1$-hop synthesis while $X^{(2)}=\{X^{(2)}_{1},X^{(2)}_{2},...,X^{(2)}_{m_{2}}\}$ represents synthetic data from $2$-hop synthesis. $r$ is the relation between a topic $t$ and a knowledge attribute $a$. The predefined operation $Op$ is an abbreviation for operation. The green area highlights a synthesis path that illustrates the relevance between two data points. The orange area shows a path to synthesize data with diversity and relevance. We zoom in one of the branches related to $X^{(1)}_{3}$ in $2$-hop synthesis. Table 1 demonstrates an example of the synthesis.
@@ 3.2.1 Multi-Hop Synthesis Guided by Attributes for Relevance
We show the prompt of extracting ${\mathcal{A}}^{(K)}_{i}$ from $X^{(K)}_{i}$ in Appendix B. Using $X^{(K-1)}_{i}$ and a triplet $\langle t,r,a\rangle^{(K-1)}_{i}$ from ${\mathcal{A}}^{(K-1)}_{i}$ based on Eq. (1), a LLM synthesizes $X^{(K)}_{i}$ with task demonstrations ${\mathcal{D}}_{T}$. The task demonstrations ${\mathcal{D}}_{T}$ includes task-related examples to guide the process of synthesis. To improve data complexity, we apply operations $Op$ (i.e., adding constraints, reasoning, and concreteness) during synthesis to enhance the quality of synthetic data Xu et al. (2024). This process is summarized as:
@@
Prompts for the synthesis process are shown in Appendix C. A multi-hop synthesis example is demonstrated in Figure 3 in Appendix LABEL:Unfolded_Multi-Hop and Table 1.
@@ 3.2.2 Multi-Hop Synthesis Guided by Personas for Diversity
Song et al. (2024) shos that fine-tuning LLMs with diverse data improves performance. However, generating diverse data at scale by LLMs requires varied prompts (Chan et al., 2024). To address this, we leverage Persona Hub (Chan et al., 2024) to enhance synthetic data diversity. For each data point, we retrieve the top-$P$ personas by using cosine similarity between its topic embedding and personas embeddings. The retrieved personas $p_{i}\in P$ guide multi-hop synthesis paths. Given a persona $p_{i}$, a data point $X^{(K-1)}_{i}$, task demonstrations ${\mathcal{D}}_{T}$, and a predefined operation $Op$, we synthesize $X^{(K)}_{i}$ as,
@@
Prompts for persona-guided synthesis are shown in Appendix D. Combining multi-hop synthesis with attributes and personas increases the volume of diverse, task-relevant synthetic data.
3.3 Residual Connection Mechanism
Multi-hop synthesis guided by attributes and personas generates diverse, relevant data, but relevance decreases as hop depth $K$ increases. For instance, synthesizing $10$-hop neighbors introduces unrelated themes (Figure 14 in Appendix E). To address this drift from the original input at deeper synthesis depths, we introduce a residual connection between the input data and its $(K-1)$-hop neighbors when $K$ exceeds a threshold $L$, as represented below,
$X^{(K)}_{i}=$ =$=\begin{cases}\text{LLM}(X^{(K-1)}_{i},\langle t,r,a\rangle^{(K-1)}_{i},Op,{\mathcal{D}}_{T}),&K<L\\ \text{LLM}(X^{(K-1)}_{i},\langle t,r,a\rangle^{(K-1)}_{i},Op,{\mathcal{D}}_{T},X^{(0)}_{i}),&K\geq L\\ \end{cases}$
Figure 15 in Appendix F demonstrates a $10$-hop synthesis using residual connections. Compared to the irrelevant content in the $10$-hop neighbor of Figure 14, the $10$-hop neighbor in Figure 15 remains focused on the relevant topic.
3.4 Self-Reflection for Quality Improvement
To ensure the correctness, relevance, and diversity of synthetic data, we adapt the self-reflection technique (Madaan et al., 2023; Pan et al., 2024) after synthesis (Figure 1). A LLM grades synthetic data $X^{(K)}_{i}$ on these aspects, providing a score (from $1$ to $10$) and feedback. Data exceeding a score threshold (i.e., threshold equal to $5$) is added to the dataset; otherwise, it undergoes limited re-synthesis iterations. A LLM annotator then labels the data, with self-reflection ensuring labeling correctness. Related prompts are shown in Appendix G.
@@ 4 Experiment
Base Model Fine-tuning with Data Source MMLU ARC- Challenge GSM8K TruthfulQA Avg. Avg. $\Delta$ ($\uparrow$)
Bio. CS Phi. EE Market.
Mistral-7B AIDE (Ours) 75.5% 57.0% 72.2% 60.7% 89.3% 74.7% 59.1% 69.2% 69.7% 23.4%
Evol-Instruct 47.1% 52.3% 68.9% 41.4% 76.1% 43.3% 60.9% 24.5% 51.8% 5.5%
Gold training data 73.2% 56.0% 71.1% 60.0% 85.9% 79.4% 53.4% 49.9% 66.1% 19.8%
Base Model 64.5% 21.0% 65.6% 35.1% 73.9% 47.5% 51.7% 11.0% 46.3% NA
Llama-3.1-8B AIDE (Ours) 74.2% 47.0% 63.0% 49.7% 82.1% 69.8% 65.8% 69.2% 65.1% 10.8%
Base Model 54.8% 43.0% 56.0% 49.0% 67.1% 66.8% 63.7% 33.9% 54.3% NA
Llama-3.2-3B AIDE (Ours) 58.7% 43.4% 56.6% 54.5% 71.4% 56.8% 45.1% 67.6% 56.8% 10.8%
Base Model 53.9% 41.2% 49.2% 49.0% 60.7% 54.3% 37.2% 22.4% 46.0% NA
Table 2: We compare the performance of different zero-shot learning methods across $5$ tasks from MMLU, ARC-Challenge, GSM8K (8-shot with maj@8) and TruthfulQA. We categorize each method by its base model and data source. We highlight the best and runner-up performance. The Avg. shows the average performance across all benchmarks. For GSM8K, we use $3.2$K gold training data to fine-tune the models, which matches the amount of synthetic data from AIDE. In contrast, Evol-Instruct uses $250$K synthetic data to fine-tune Mistral-7B. We obtain the results using the same parameter settings, without performing hyperparameters tuning. NA indicates that there is no difference between the base model and itself.
In this section, we evaluate AIDE to answer the following research questions (RQs):
  • (RQ1) Can AIDE improve base models’ performance through fine-tuning?
  • (RQ2) How does AIDE affect base models’ performance under different settings?
  • (RQ3) Does the data from AIDE maintain relevance, diversity and complexity?

### v2

# AIDE: Attribute-Guided MultI-Hop Data Expansion for Data Scarcity in Task-Specific Fine-tuning

@@ Abstract
Fine-tuning large language models (LLMs) for specific tasks requires diverse, high-quality training data. However, obtaining sufficient relevant data remains a significant challenge. Existing data synthesis methods either depend on extensive seed datasets or struggle to balance task relevance and data diversity. To address these challenges, we propose Attribute-guided multI-hop Data Expansion (AIDE), a novel data synthesis framework that uses a multi-hop process to expand very few seed data points while ensuring data diversity and task relevance. AIDE extracts the main topic and key knowledge attributes from the seeds to guide the synthesis steps. The process repeats for $K$ hops, using the generated data as seeds. To prevent irrelevant data generation as the hop depth increases, AIDE incorporates a residual connection mechanism. Our empirical results show that AIDE enables fine-tuning of Mistral-7B, Llama-3.1-8B and Llama-3.2-3B from 10 seeds, surpassing the models fine-tuned on human curated data. Furthermore, AIDE outperforms state-of-the-art data synthesis methods, such as Evol-Instruct, by over $30\%$ in task-specific fine-tuning. Code is available at https://github.com/Code4Graph/AIDE.
@@
AIDE: Attribute-Guided MultI-Hop Data Expansion
for Data Scarcity in Task-Specific Fine-tuning
Jiayu Li1,3,thanks: Work was done during an internship at AWS., Xuan Zhu2, Fang Liu2, Yanjun Qi2,3 1Syracuse University, 2AWS Bedrock Science 3Correspondence: jli221@data.syr.edu, yanjunqi@amazon.com
@@ 1 Introduction
To overcome the limitation, an approach from Wei et al. (2022); Xu et al. (2022) samples task-specific training data from public NLP datasets, but the sampling often covers limited information. Another category of recent methods leverages the capabilities of LLMs to automatically generate large-scale synthetic data, enabling the training of advanced models in specific task domains. For example, Prompt2Model (Viswanathan et al., 2023) and DataTune (Gandhi et al., 2024) rely on several candidate datasets to synthesize task-specific data for fine-tuning LLMs. However, these methods either require a large set of seed data for rewriting or produce synthetic data that lacks task relevance and diversity, as they do not maintain sufficient control over the synthesis process.
@@
To address these challenges, we propose AIDE (Attribute-guided multI-hop Data Expansion), a novel data synthesis framework that generates abundant training data from a small set of seed inputs, as shown in Figure 1. Our framework focuses on maintaining high task relevance, diversity, and quality in the synthetic data for specific tasks. AIDE uses LLMs as key players via a multi-hop synthesis process. Each hop in AIDE begins by extracting the main topic and important knowledge attributes from a seed sample using a LLM. This builds knowledge triplets, and AIDE traverses these triplets (each consisting of a topic, relationship, and attribute) to synthesize new data points. In the next hop, each newly generated data point becomes a seed, and the process repeats until reaching a depth of $K$ hops. This multi-hop mechanism allows for recursive data synthesis along all paths of a process tree, enabling the generation of large-volume data from just a few seeds. Extracted attributes act as control nodes in the multi-hop tree, ensuring the generated data points remain relevant to the target task. We also introduce personas as new key attributes, enhancing the generation of diverse data. As the depth of the recursive synthesis increases, the relevance of the synthetic data may diminish. To address this, we propose a residual connection mechanism to reduce irrelevance.
@@
To validate AIDE, we conduct experiments with three pretrained models (Mistral-7B, Llama-3.1-8B, and Llama-3.2-3B). We evaluate the performance of these models when fine-tuned with synthetic data generated by AIDE, comparing the results against models fine-tuned with human-curated (gold) data and synthetic data from state-of-the-art (SOTA) methods. Our evaluations span a range of tasks from well-known benchmarks, including industrial datasets like MedQA Jin et al. (2020) and FinBen Xie et al. (2024), as well as BIG-Bench bench authors (2023), MMLU Hendrycks et al. (2021), ARC-Challenge Clark et al. (2018), GSM8K Cobbe et al. (2021), and TruthfulQA Lin et al. (2022). For comparison, we include SOTA data synthesis methods such as Evol-Instruct Xu et al. (2024), DataTune Gandhi et al. (2024), and Prompt2Model Viswanathan et al. (2023). Our main contributions are as follows:
@@
  • In zero-shot prompting, Mistral-7B fine-tuned with synthetic data from AIDE achieves average relative improvements of over $6\%$ and $30\%$ across tasks, compared to Mistral-7B fine-tuned with gold training data and SOTA data synthesis methods. Additionally, AIDE enhances the performance of Llama-3.1-8B and Llama-3.2-3B, yielding average relative improvements of approximately $0.7\%$ and $1.5\%$ across tasks, respectively, compared to fine-tuning with gold data.
@@ 2 Related Work
Data synthesis for fine-tuning LLMs targets two primary problems. The first is open-domain generation, which synthesizes data across a wide range of topics and complexity levels. The second is task-specific generation, where synthetic data is tailored to a particular task. One can use the synthetic data in fine-tuning LLMs through techniques, such as instruction tuning, preference tuning, and their variations. This paper focuses on synthesizing training data for instruction tuning to enhance the performance of LLMs for specific tasks. We discuss related methods for data synthesis in both open and task-specific domains in Appendix A.
@@
3 Proposed Method: Attribute-Guided Multi-Hop Data Expansion (AIDE)
Figure 1: Overview of the workflow of AIDE. $X^{(0)}_{i}$ denotes the $i$-th task-related seed data point. AIDE includes four steps. (1) a LLM extractor extracts a topic $t$, knowledge attributes $a_{1}$ and $a_{2}$ with relationships $r_{1}$ and $r_{2}$ of a data point. (2) During the multi-hop synthesis at the depth of hop $j$, a LLM acts as a synthesizer with task demonstrations ${\mathcal{D}}_{T}$ to generate data $X^{(j)}_{1}$, $X^{(j)}_{2}$ and $X^{(j)}_{3}$ along paths of synthesis with a predefined operation $Op$ (i.e., adding constraints). (3) To enhance the diversity of synthesis, we expand attributes by retrieving a persona $p_{1}$ from a persona hub with $t$. Finally, a LLM as an annotator generates the label of synthetic data. We describe the technical details of AIDE in Section 3.
@@ 3 Proposed Method: Attribute-Guided Multi-Hop Data Expansion (AIDE)
Table 1: The $1$-hop synthesis in Figure 9 of Appendix C uses an input data point $X^{(0)}_{i}$ to generate a representation of the data point ${\mathcal{A}}^{(0)}_{i}$ with triplets $<t_{1},r_{1},a_{1}>$ and $<t_{1},r_{2},a_{2}>$. We retrieve the persona $P_{1}$ according to $t_{1}$. Through the triplets, task demonstrations ${\mathcal{D}}_{T}$, the persona $p_{1}$ and the predefined operation $Op$, we synthesize $X^{(1)}_{1}$, $X^{(1)}_{2}$ and $X^{(1)}_{3}$ by combining $X^{(0)}_{i}$ with its corresponding task category and related examples.
3.1 Multi-Hop Synthesis
To synthesize abundant data, we propose a multi-hop synthesis approach, with an example illustrated in Figure 8 of Appendix B.
Definition 3.1 (Multi-hop synthesis).
@@ 3.2.1 Multi-Hop Synthesis Guided by Attributes for Relevance
We show the prompt of how to extract ${\mathcal{A}}^{(K)}_{i}$ from $X^{(K)}_{i}$ in Appendix I. Using $X^{(K-1)}_{i}$ and a triplet $\langle t,r,a\rangle^{(K-1)}_{i}$ from ${\mathcal{A}}^{(K-1)}_{i}$ based on Eq. (1), a LLM synthesizes $X^{(K)}_{i}$ with task demonstrations ${\mathcal{D}}_{T}$. The task demonstrations ${\mathcal{D}}_{T}$ includes task-related examples to guide the process of synthesis. To improve data complexity, we apply operations $Op$ (i.e., adding constraints, reasoning, and concreteness) during synthesis to enhance the quality of synthetic data Xu et al. (2024). This process is summarized as:
@@
Prompts for the synthesis process are shown in Appendix J. A multi-hop synthesis example is demonstrated in Figure 9 in Appendix C and Table 1.
@@ 3.2.2 Multi-Hop Synthesis Guided by Personas for Diversity
Song et al. (2024) shows that fine-tuning LLMs with diverse data improves performance. However, generating diverse data at scale by LLMs requires varied prompts (Chan et al., 2024). To address this, we leverage Persona Hub (Chan et al., 2024) to enhance synthetic data diversity. For each data point, we retrieve the top-$P$ personas by using cosine similarity between its topic embedding and personas embeddings. The retrieved personas $p_{i}\in P$ guide multi-hop synthesis paths. Given a persona $p_{i}$, a data point $X^{(K-1)}_{i}$, task demonstrations ${\mathcal{D}}_{T}$, and a predefined operation $Op$, we synthesize $X^{(K)}_{i}$ as,
@@
Prompts for persona-guided synthesis are shown in Appendix K. Combining multi-hop synthesis with attributes and personas increases the volume of diverse, task-relevant synthetic data.
3.3 Residual Connection Mechanism for Maintaining Task Relevance
Multi-hop synthesis guided by attributes and personas generates diverse, relevant data, but relevance decreases as hop depth $K$ increases. For instance, synthesizing $10$-hop neighbors introduces unrelated themes (Figure 15 in Appendix L). To address this drift from the original input at deeper synthesis depths, we introduce residual connections between a seed data point and its neighbors. Specifically, for any depth $d$ where $1<d\leq K$, we build the connections when $d\leq L$ where $L$ is the depth of residual connection within the range $(1,K]$,
$X^{(d)}_{i}=$$\begin{cases}\text{LLM}(X^{(d-1)}_{i},\langle t,r,a\rangle^{(d-1)}_{i},Op,{\mathcal{D}}_{T}),&L<d\\ \text{LLM}(X^{(d-1)}_{i},\langle t,r,a\rangle^{(d-1)}_{i},Op,{\mathcal{D}}_{T},X^{(0)}_{i}),&d\leq L.\\ \end{cases}$
We illustrate the detail of residual connection in Appendix D. Figure 16 demonstrates a $10$-hop synthesis using residual connections. Compared to Figure 15, the $10$-hop neighbor in Figure 16 remains focused on the relevant topic.
@@ 4 Experiment
Fine-tuning with Data Source MMLU FinBen ARC- Challenge GSM8K TruthfulQA MedQA Avg. ($\uparrow$) Avg. $\Delta$ ($\uparrow$)
Bio. CS Phi. EE Market. CFA
# Seed Data Points in AIDE 10 10 10 10 10 10 10 10 10 10 10
Pretrained Mistral-7B AIDE (Ours) 75.5% 57.0% 72.2% 60.7% 89.3% 41.0% 74.7% 59.1% 69.2% 44.0% 64.3% 7.0%
Gold training data 73.2% 56.0% 71.1% 60.0% 85.9% 35.0% 79.4% 53.4% 49.9% 37.0% 60.1% NA
Pretrained Llama-3.1-8B AIDE (Ours) 74.2% 47.0% 63.0% 49.7% 82.1% 62.0% 69.8% 65.8% 69.2% 56.0% 63.9% 0.7%
Gold training data 74.7% 48.1% 60.5% 50.1% 82.3% 61.0% 69.6% 68.2% 66.1% 54.0% 63.7% NA
Pretrained Llama-3.2-3B AIDE (Ours) 58.7% 43.4% 56.6% 54.5% 71.4% 54.0% 56.8% 45.1% 67.6% 51.0% 55.9% 1.5%
Gold training data 60.2% 45.0% 55.6% 48.3% 70.7% 54.0% 56.5% 45.5% 64.9% 50.0% 55.1% NA
Table 2: AIDE-generated data vs. human-curated training data for fine-tuning. We evaluate the performance of various zero-shot learning methods across MMLU, FinBen, ARC-Challenge, GSM8K (8-shot with maj@8), TruthfulQA, and MedQA. We highlight the best and runner-up performances. "Avg." represents the average performance across all benchmarks. For GSM8K, we fine-tune the models using 3.2K gold training data, matching the amount of synthetic data from AIDE. Results are obtained using the same parameter settings. Avg. $\Delta(\uparrow)$ represents the relative average improvement of models compared to those fine-tuned with gold data. "NA" indicates no difference from models fine-tuned with gold data.
We evaluate AIDE to answer the following research questions (RQs): (RQ1) Can AIDE enable the fine-tuning of pretrained models that outperform those fine-tuned on human-curated data and data generated by SOTA synthesis methods? (RQ2) How does AIDE affect pretrained models’ performance under different settings? (RQ3) Does the data from AIDE maintain relevance and diversity?